# Invariance transformations in wave-optics lensing: implications for gravitational-wave astrophysics and cosmology

### v1

@@ VII Conclusions
As for the cosmology studies, we analyzed the degeneracy caused by the change of the scaled Hubble constant relative to the fiducial value, $h_{0}$, which is degenerate with the lens mass due to the invariance transformation, and with the luminosity distance of the source. We then performed a population study by simulating lensed GW events following a LVK source mass distribution model, and assuming a lens density uniform in comoving volume and fixed lens mass of $100\,M_{\odot}$. We used the Fisher matrix to forecast the constraints on $h_{0}$ at LIGO A+ sensitivity and ET sensitivity. While the uncertainties in $h_{0}$ are over 100% for most of the events with current sensitivity, many of those with next-generation sensitivity (such as ET) reduce to the order of $10\%$. Better constraints could be achieved for larger lens masses. Our forecast reveals a possibility to measure the Hubble constant with wave-optics lensed GW events. A more accurate forecast needs to be done with a better lens population model and it will be worth also exploring future space-based detectors like LISA Amaro-Seoane et al. (2017) that could detect very high SNR events.
@@ Appendix A The second minimum in mismatch
As shown in the lower panel of Figure 1, the mismatch computed by aligning the waveforms with the fiducial time delay has a second minimum apart from the fiducial point of $M_{L}=700,\lambda=1$. It is located along the curve of a constant product $M_{L}\lambda$, at a lower $\lambda$ value. We plotted the mismatch along this constant product curve in one dimension to examine the second minimum for the PM model, as shown in Figure 14. Unlike the minimum at the fiducial value, the second minimum does not drop to 0. We further plotted the mismatch for different values of $y$, and found that the second minimum point varies with $y$. However, it does not vary with any waveform parameters. Moreover, when we plotted the one-dimensional mismatch along the constant product curve of $M_{L}\lambda$ for the SIS model, there is no second minimum, for different values of $y$. Therefore this feature of a second minimum is specific to the PM lens model. This feature is also present in Fig. 7 of Cremonese et al. (2021) as a second peak of the ratio between SNR and the optimal SNR.
@@ Appendix B Parameter estimation analyses
The injections are done using the pycbc package Nitz et al. (2024). The PSD used is the aLIGOZeroDetHighPower555dcc.ligo.org/LIGO-T070247/public.. For the injections, we assume detection at three detectors: Hanford, Livingston Aasi et al. (2015) and Virgo Acernese et al. (2014). The amplification factor has been computed in python and Mathematica in the form of a look-up table, following Eqs. (33) and (41) for PM and SIS lens, respectively. The PE runs have been done using bilby 2.2.3 through bilby-pipe 1.3.0 Ashton et al. (2019).
@@ B.1 PM injection
For this first case, we will use an injection described by the parameters reported in Table 1. We choose the source parameters to have an event with a high SNR and the lens parameters such that we are in the wave-optics regime in the relevant frequency band666 The condition for which the geometrical optics approximation is broken is given by Takahashi (2017) $M_{3D,L}\leq 10^{5}M_{\odot}\left[1+z_{L})f/\text{Hz}\right]^{-1}$, where $M_{3D,L}$ and $z_{L}$ are the 3D mass and redshift of the lens, respectively. This condition is, in general, dependent on $y$, and holds true for $y\sim 1$ Ezquiaga et al. (2021); Bulashenko and Ubach (2022). .
@@
As we can see, the uncertainties on the lens parameters are rather low, i.e. $\approx 4\%$, and in line with previous results Cremonese et al. (2021). As we shall see, though, these are not the “real” uncertainties, since we are not considering the MSD yet. All the parameters are retrieved correctly without any bias. This is to be expected since the injection with $\lambda=1$ is included in the PE model without $\lambda$, i.e. $\lambda=1$ is the case where no transformation is present and, therefore, we are not yet considering any source of degeneracy.
@@ Appendix C Mass distribution and lensing rate in population study
In this appendix, we present the plots showing the detailed intermediate steps in the population study for wave-optics cosmography discussed in Sec. VI.2. Figure 23 shows the Powerlaw + peak mass distribution model generated with parameters listed in Sec. VI.2, using the GWCOSMO package Gray et al. (2020, 2022, 2023). The component masses of the simulated binary merger events are randomly drawn from this probability distribution, with $m_{1}>m_{2}$.

### v2

@@ VII Conclusions
As for the cosmology studies, we analyzed the degeneracy caused by the change of the scaled Hubble constant relative to the fiducial value, $h_{0}$, which is degenerate with the lens mass due to the invariance transformation, and with the luminosity distance of the source. We then performed a population study by simulating lensed GW events following a LVK source mass distribution model, and assuming a lens density uniform in comoving volume and fixed lens mass of $100\,M_{\odot}$. We used the Fisher matrix to forecast the constraints on $h_{0}$ at LIGO A+ sensitivity and ET sensitivity. While the uncertainties in $h_{0}$ are over 100% for most of the events with current sensitivity, many of those with next-generation sensitivity (such as ET) reduce to the order of $10\%$. Better constraints could be achieved for larger lens masses. Our forecast reveals a possibility to measure the Hubble constant with wave-optics lensed GW events. A more accurate forecast needs to be done with a better lens population model and it will be worth also exploring future space-based detectors like LISA [113] that could detect very high SNR events.
@@ Appendix A The second minimum in mismatch
As shown in the lower panel of Figure 1, the mismatch computed by aligning the waveforms with the fiducial time delay has a second minimum apart from the fiducial point of $M_{L}=700,\lambda=1$. It is located along the curve of a constant product $M_{L}\lambda$, at a lower $\lambda$ value. We plotted the mismatch along this constant product curve in one dimension to examine the second minimum for the PM model, as shown in Figure 14. Unlike the minimum at the fiducial value, the second minimum does not drop to 0. We further plotted the mismatch for different values of $y$, and found that the second minimum point varies with $y$. However, it does not vary with any waveform parameters. Moreover, when we plotted the one-dimensional mismatch along the constant product curve of $M_{L}\lambda$ for the SIS model, there is no second minimum, for different values of $y$. Therefore this feature of a second minimum is specific to the PM lens model. This feature is also present in Fig. 7 of [71] as a second peak of the ratio between SNR and the optimal SNR.
@@ Appendix B Parameter estimation analyses
The injections are done using the pycbc package [114]. The PSD used is the aLIGOZeroDetHighPower555dcc.ligo.org/LIGO-T070247/public.. For the injections, we assume detection at three detectors: Hanford, Livingston [1] and Virgo [2]. The amplification factor has been computed in python and Mathematica in the form of a look-up table, following Eqs. (33) and (41) for PM and SIS lens, respectively. The PE runs have been done using bilby 2.2.3 through bilby-pipe 1.3.0 [115].
@@ B.1 PM injection
For this first case, we will use an injection described by the parameters reported in Table 1. We choose the source parameters to have an event with a high SNR and the lens parameters such that we are in the wave-optics regime in the relevant frequency band666 The condition for which the geometrical optics approximation is broken is given by [65] $M_{3D,L}\leq 10^{5}M_{\odot}\left[1+z_{L})f/\text{Hz}\right]^{-1}$, where $M_{3D,L}$ and $z_{L}$ are the 3D mass and redshift of the lens, respectively. This condition is, in general, dependent on $y$, and holds true for $y\sim 1$ [10, 116]. .
@@
As we can see, the uncertainties on the lens parameters are rather low, i.e. $\approx 4\%$, and in line with previous results [71]. As we shall see, though, these are not the “real” uncertainties, since we are not considering the MSD yet. All the parameters are retrieved correctly without any bias. This is to be expected since the injection with $\lambda=1$ is included in the PE model without $\lambda$, i.e. $\lambda=1$ is the case where no transformation is present and, therefore, we are not yet considering any source of degeneracy.
@@ Appendix C Mass distribution and lensing rate in population study
In this appendix, we present the plots showing the detailed intermediate steps in the population study for wave-optics cosmography discussed in Sec. VI.2. Figure 23 shows the Powerlaw + peak mass distribution model generated with parameters listed in Sec. VI.2, using the GWCOSMO package [98, 99, 100]. The component masses of the simulated binary merger events are randomly drawn from this probability distribution, with $m_{1}>m_{2}$.